Oldham Inc. conducts business in State M and State N, which both use the UDITPA three-factor formula to apportion income. State M’s corporate tax rate is 4.50 percent, and State N’s corporate tax rate is 7.00 percent. This year, Oldham had the following sales, payroll, and property (in thousands of dollars) in each state:

State M State N Total
Gross receipts from sales $ 3,600 $ 8,100 $ 11,700
Payroll expense 1,400 1,800 3,200
Property costs 1,500 1,600 3,100

If Oldham’s before-tax income was $3.6 million, compute its State M and State N tax. (Round all apportionment factors to 4 decimal places, e.g., 47.15% would be 0.4715. Enter "Apportionment percentage" and "Tax rate" in decimals and not in percentage. Enter your answers in dollars and not in millions or thousands of dollars.) STATE M STATE N

Total taxable income subject to apportionment:   

Apportionment Percentage:

State taxable income:

Tax Rate:

State Tax:
Answer #1

A B 1 State M 2 Sales factor 3 payroll factor 4 property factor 5 Total Apportionment percentage State N 30.77 43.75 48.39 12C A 1 2 Sales factor 3 payroll factor 4 property factor 5 Total B State M State N =(3600/11700)*100 =(8100/11700)*100 =(1400/